The probability that a certain NBA player makes a free throw is 77%. In one game, he attempted 12 free throws.
A. What is the probability that he made exactly 3 of them?

B. What is the probability that he made all but one of them?

C. What is the probability that he made more than half of them?

Answer by Boreal(15235)

Exactly 3--2 VARS binomial pdf ENTER(12,.77,3)=0.00018
all but one of them is 12*.77^11*.23=0.1557
for more than half, do 1- half or less or 1-binomialCdf(12..77,6)=0.9626
mean is 9.24 (np)
variance is np(1-p)=9.24*0.23=2.13
sd is sqrt (V)=1.46
This is a crude check, but using continuity correction factor z=(6.5-9.24)/1.46
=-1.876
probability z >-1.876 is 0.9697